A population of pea plants may have white flowers (the dominant phenotype) or yellow flowers (the recessive phenotype). The frequency of homozygous dominant individuals is 0.63. Answer the questions below. For each, type out the equation with the appropriate numbers filled in as well as the answer.

How would you calculate the frequency of the white homozygous alleles?
How would you calculate the frequency of the recessive alleles?
How would you calculate the frequency of the heterozygous individuals?
Answer #1

Answer - According to Hardy Weinberg Law :

P2 + Q2 + 2PQ = 1

here P is frequency of dominant allele responsible for white flowers

and Q is the frequency of recessive allele responsible for yellow flowers

Answer 1

Given in question that the frequency of homozygous dominant individuals is 0.63

means >> P2 = 0.63

means >> P= (0.63)1/2

means >> P= 0.79

so, Answer 1. Frequency of White homozygous alleles is 0.79

Answer 2

we also know that >>   P + Q = 1

so, Q  = 1 - 0.79

means Q = 0.21

so Answer 2 Frequency of recessive alleles is 0.21

Answer 3

According to Hardy Weinberg Law :

P2 + Q2 + 2PQ = 1 ( here 2PQ is nothing but frequency of heterozygous individuals)

so putting values of P2 and Q2 in above equation we get :

0.63 + 0.04 + 2PQ = 1

means 2PQ = 1 - 0.63 - 0.04

means 2PQ = 0.33

So, Answer 3. Frequency of heterozygous individuals is 0.33